# Restriction of eigenfunctions on products of spheres to submanifolds of maximal flats

### v1

@@ 1. Introduction
Let $M$ be a compact Riemannian manifold of dimension $d$. Let $\Delta$ be the Laplace–Beltrami operator on $M$, and let $f$ be its eigenfunctions such that $\Delta f=-N^{2}f$, $N>1$. The study of concentration of $f$ has become a popular subject [Zel17], and one of the approaches is to bound its restriction to submanifolds of $M$, as initiated by Tataru [Tat98] and Reznikov [Rez04]. Let $S$ be a submanifold of dimension $k$. Let $L^{p}(M)$, $L^{p}(S)$ be the Lebesgue spaces associated to the Riemannian volume measure on $M$ and $S$ respectively. The following fundamental restriction bounds for general $M$ were established by Burq–Gérard–Tzvetkov [BGT07] and Hu [Hu09]: it holds
@@
All the above estimates are sharp on spheres, except for the logarithmic loss, which however has been eliminated by Chen–Sogge [CS14] and Wang–Zhang [WZ21] for totally geodesic submanifolds.
@@
In the setting of compact Lie groups, the author established the same sharp estimate as (1.4) for eigenfunctions restricted to submanifolds of maximal flats [Zha26].
@@
In the setting of standard tori, Huang–Zhang [HZ21] proved that for totally geodesic submanifolds defined by linear equations with rational coefficients, the same eigenfunction restriction estimate as (1.4) holds for all $p\geq 2$, up to an $\varepsilon$-loss.
@@
In particular, our estimates apply to geodesics, since every geodesic lies in a maximal flat of $M$. It remains an interesting question to get sharp estimates for restriction of eigenfunctions to submanifolds which are not contained in any maximal flat. A possible approach, which we leave for future work, is to adapt the argument of Burq–Gérard–Tzvetkov [BGT07] to the product manifold setting.
@@
For global $L^{p}$ estimates of eigenfunctions, the author established on general symmetric spaces of compact type that [Zha21]
@@
for $r\geq 5$ and p>2+8/(r4)p>2+{8}/{(r-4)}. The above sharp estimate also holds on products of CROSSs for $r\geq 5$ and $p\geq{2(d_{i}+1)}/{(d_{i}-1)}$ for all $i$, by applying Sogge’s classical spectral projector estimates [Sog88] to each factor. It remains open to establish sharp $L^{p}$ estimates for all $p\geq 2$.
Our approach to proving Theorem 1.1 is to extract information about general eigenfunctions from a special class of eigenfunctions, namely the spherical functions. A key ingredient is the mapping property of convolution with Jacobi polynomials, which express spherical functions on CROSSs. Let $P_{n}^{(\alpha,\beta)}(x)$ ($n\in\mathbb{Z}_{\geq 0}$) denote the Jacobi polynomials, defined as the orthogonal polynomials on the interval [1,1][-1,1] with the weight function $(1-x)^{\alpha}(1+x)^{\beta}$. We follow the classical treatment in Szegö’s book [Sze75]. In particular, we use the same normalization
@@
In particular, the case $\alpha=\beta=\tfrac{1}{2}$ corresponds to the three-dimensional factors $\mathbb{S}^{3}$ or $\mathbb{RP}^{3}$ in the product manifold $M$. For this case, the sharp estimate (1.8) at the endpoint $p=2$ mirrors the log-loss removal of (1.2) for geodesics on three-dimensional compact manifolds established by Chen–Sogge [CS14]. In our setting of product manifolds containing three-dimensional factors, another key ingredient that enables the use of (1.8) for eigenfunction restriction to arbitrary submanifolds of maximal flats is the positivity of Fourier coefficients of spherical functions, as in Lemma 2.1. This positivity helps preserve oscillation in the convolution kernel and effectively reduces the study of arbitrary submanifolds to those of product type; see (2.9).
@@ Acknowledgments
I would like to thank Xiaocheng Li and Cheng Zhang for helpful discussions.
@@ 2.1. Spherical functions
In particular, we have picked the standard normalization such that $\Phi_{n}(0)=1$. For the above formula, we refer to Theorem 4.5 in Chapter V of [Hel00], where the $\Phi_{n}$ are expressed in terms of hypergeometric functions, and to formula (4.21.2) in [Sze75], where these hypergeometric functions are rewritten as Jacobi polynomials. A specific case is when $d=3\Leftrightarrow\alpha=1/2$, corresponding to $\mathbb{S}^{3}$ or $\mathbb{RP}^{3}$, for which we also have $\beta=1/2$.
@@
Lemma 2.1 (Proposition 9.4 of Chapter III of [Hel08]).
@@
This is a consequence of the Weyl dimension formula combined with the characterization of the highest weight of spherical representations (Theorem 4.1 of Chapter V of [Hel00]). ∎
@@ 2.2. Eigenfunctions
For $f\in L^{2}(M)$, following Proposition 9.1 of Chapter III of [Hel00], we have the $L^{2}$ decomposition
@@
where $\mathbb{T}_{i}$ is a maximal torus of $M_{i}$. By taking a finite open cover of $S$, we may assume there exist $1\leq i_{1}<\cdots<i_{k}\leq r$ such that $S$ is parametrized by the map
$\mathbb{T}^{k}\supset V\ni\underline{\theta}=(\theta_{i_{1}},\ldots,\theta_{i_{k}})\mapsto(\theta_{1}(\underline{\theta}),\ldots,\theta_{r}(\underline{\theta}))\in S\subset E.$
@@
$=\prod_{i=1}^{r}k_{i}(n_{i})\cdot\prod_{l=1}^{k}\binom{n+\alpha_{i_{l}}}{n}^{-1}P_{n_{i_{l}}}^{(\alpha_{i_{l}},\beta_{i_{l}})}(\cos(\theta_{i_{l}}-\theta_{i_{l}}^{\prime})),$
@@
(2.10) $\mathcal{A}:g\mapsto\int_{V}g(\underline{\theta}^{\prime})\widetilde{\mathscr{K}}(\underline{\theta}^{\prime},\underline{\theta})\ d\underline{\theta}^{\prime}.$
Furthermore, we may now assume $V=\mathbb{T}^{k}$ without loss of generality, and then the above operator is convolution on $\mathbb{T}^{k}$ with a product of Jacobi polynomials. We may apply Theorem 1.8 along with the Minkowski inequality (using $p\geq 2\Leftrightarrow p\geq p^{\prime}$) to bound such a convolution operator, as well as apply the standard asymptotics
@@
and Lemma 2.2 to bound the coefficients, to obtain
$\|\mathcal{A}\|_{L^{p^{\prime}}(S)\to L^{p}(S)}\lesssim N^{{d-r}+2\sum_{l=1}^{k}\tau(d_{i_{l}},p)}\leq N^{{d-r}+2\sum_{i=1}^{k}\tau(d_{i},p)}.$
@@
Suppose $\Delta f=-N^{2}f$, $N>1$. By (2.4), and a standard counting estimate [Gro85], we get
@@
We decompose the convolution kernel $P^{\alpha,\beta}_{n}(\cos\theta)$ according to the value of $\theta$, and it suffices to restrict attention to $\theta\in[0,\pi]$. Let $c$ be a fixed positive constant. Let $\widetilde{n}:=n+(\alpha+\beta+1)/2$, $\gamma:=-(\alpha+\frac{1}{2})\pi/2$.
Case 1. $c(n+1)^{-1}\leq\theta\leq\pi-c(n+1)^{-1}$. We use (8.21.18) of [Sze75] as follows:
@@
Case 2. $0<\theta\leq c(n+1)^{-1}$ or $\pi-c(n+1)^{-1}\leq\theta<\pi$. For the former, we use (8.21.17) of [Sze75] as follows:
@@
Here $J_{\alpha}$ stands for the Bessel function of the first kind of order $\alpha$, as defined in (1.71.1) of [Sze75]. By (1.71.10) of [Sze75], we have the asymptotics,
@@
A similar approach works for the other case $\pi-c(n+1)^{-1}\leq\theta<\pi$: combining (8.21.17) and (4.1.3) of [Sze75], we get
@@
at the kink point $p={4}/{(2\alpha+1)}>2$. To accomplish this, we adjust the proof of Proposition 3.1: the only change needed is to apply the Hardy–Littlewood–Sobolev inequality in place of Young’s convolution inequality to (3.1) over the interval $c(n+1)^{-1},\pi/2]$, as well as over the interval $\pi/2,\pi-c(n+1)^{-1}]$ if $\beta=\alpha$; if $\beta<\alpha$, the original application of Young’s inequality over the interval $\pi/2,\pi-c(n+1)^{-1}]$ via (3.2) remains sufficient. For the remaining case $\alpha=\beta=1/2$, we use the explicit formula (4.1.7) of [Sze75],
@@
As the spherical functions $\Phi_{i,n_{i}}$ are matrix entries of the associated irreducible spherical representations expressed via unit vectors (see Theorem 4.3 of Chapter V of [Hel00]), the Schur orthogonality relations yield orthogonality between $\Phi_{i,n_{i}}$ for different $n_{i}$, and
@@
When $r\geq 5$, by the equidistribution of lattice points on $r$-dimensional spheres as established by Pommerenke [Pom59], it holds $|\Lambda|\gtrsim N^{r-2}$ whenever $\Lambda\neq\varnothing$, which yields the sharpness of (1.4).

### v2

@@ 1. Introduction
Let $M$ be a compact Riemannian manifold of dimension $d$. Let $\Delta$ be the Laplace–Beltrami operator on $M$, and let $f$ be its eigenfunctions such that $\Delta f=-N^{2}f$, $N>1$. The study of concentration of $f$ has become a popular subject [14], and one of the approaches is to bound its restriction to submanifolds of $M$, as initiated by Tataru [12] and Reznikov [9]. Let $S$ be a submanifold of dimension $k$. Let $L^{p}(M)$, $L^{p}(S)$ be the Lebesgue spaces associated to the Riemannian volume measure on $M$ and $S$ respectively. The following fundamental restriction bounds for general $M$ were established by Burq–Gérard–Tzvetkov [1] and Hu [6]: it holds
@@
All the above estimates are sharp on spheres, except for the logarithmic loss, which however has been eliminated by Chen–Sogge [2] and Wang–Zhang [13] for totally geodesic submanifolds.
@@
In the setting of compact Lie groups, the author established the same sharp estimate as (1.4) for eigenfunctions restricted to submanifolds of maximal flats [16].
@@
In the setting of standard tori, Huang–Zhang proved that for totally geodesic submanifolds defined by linear equations with rational coefficients, the same eigenfunction restriction estimate as (1.4) holds for all $p\geq 2$, up to an $\varepsilon$-loss [7].
@@
In particular, our estimates apply to geodesics, since every geodesic lies in a maximal flat of $M$. It remains an interesting question to get sharp estimates for restriction of eigenfunctions to submanifolds which are not contained in any maximal flat. A possible approach, which we leave for future work, is to adapt the argument of Burq–Gérard–Tzvetkov [1] to the product manifold setting.
@@
For global $L^{p}$ estimates of eigenfunctions, the author established on general symmetric spaces of compact type that
@@
for $r\geq 5$ and p>2+8/(r4)p>2+{8}/{(r-4)} [15]. The above sharp estimate also holds on products of CROSSs for $r\geq 5$ and $p\geq{2(d_{i}+1)}/{(d_{i}-1)}$ for all $i$, by applying Sogge’s classical spectral projector estimates [10] to each factor. It remains open to establish sharp $L^{p}$ estimates for all $p\geq 2$.
Our approach to proving Theorem 1.1 is to extract information about general eigenfunctions from a special class of eigenfunctions, namely the spherical functions. A key ingredient is the mapping property of convolution with Jacobi polynomials, which express spherical functions on CROSSs. Let $P_{n}^{(\alpha,\beta)}(x)$ ($n\in\mathbb{Z}_{\geq 0}$) denote the Jacobi polynomials, defined as the orthogonal polynomials on the interval [1,1][-1,1] with the weight function $(1-x)^{\alpha}(1+x)^{\beta}$. We follow the classical treatment in Szegö’s book [11]. In particular, we use the same normalization
@@
In particular, the case $\alpha=\beta=\tfrac{1}{2}$ corresponds to the three-dimensional factors $\mathbb{S}^{3}$ or $\mathbb{RP}^{3}$ in the product manifold $M$. For this case, the sharp estimate (1.8) at the endpoint $p=2$ mirrors the log-loss removal of (1.2) for geodesics on three-dimensional compact manifolds established by Chen–Sogge [2]. In our setting of product manifolds containing three-dimensional factors, another key ingredient that enables the use of (1.8) for eigenfunction restriction to arbitrary submanifolds of maximal flats is the positivity of Fourier coefficients of spherical functions, as in Lemma 2.1. This positivity helps preserve oscillation in the convolution kernel and effectively reduces the study of arbitrary submanifolds to those of product type; see (2.9).
@@ Acknowledgments
The author thanks the referee for their valuable comments, which improved the paper. The author also thanks Xiaocheng Li and Cheng Zhang for helpful discussions.
@@ 2.1. Spherical functions
In particular, we have picked the standard normalization such that $\Phi_{n}(0)=1$. For the above formula, we refer to Theorem 4.5 in Chapter V of [4], where the $\Phi_{n}$ are expressed in terms of hypergeometric functions, and to formula (4.21.2) in [11], where these hypergeometric functions are rewritten as Jacobi polynomials. A specific case is when $d=3\Leftrightarrow\alpha=1/2$, corresponding to $\mathbb{S}^{3}$ or $\mathbb{RP}^{3}$, for which we also have $\beta=1/2$.
@@
Lemma 2.1 (Proposition 9.4 of Chapter III of [5]).
@@
This is a consequence of the Weyl dimension formula combined with the characterization of the highest weight of spherical representations (Theorem 4.1 of Chapter V of [4]). ∎
@@ 2.2. Eigenfunctions
For $f\in L^{2}(M)$, following Proposition 9.1 of Chapter III of [4], we have the $L^{2}$ decomposition
@@
where $\mathbb{T}_{i}$ is a maximal torus of $M_{i}$. Let $\theta_{1},\ldots,\theta_{r}$ denote the corresponding angular variables on $E$. By passing to a finite open cover of $S$, we may assume that there exist $1\leq i_{1}<\cdots<i_{k}\leq r$ and an open subset $V$ of $\mathbb{T}_{i_{1}}\times\cdots\times\mathbb{T}_{i_{k}}$ such that the angular variables $\theta_{i_{1}},\ldots,\theta_{i_{k}}$ among $\theta_{1},\ldots,\theta_{r}$ provide coordinates on $S$. Denote $\underline{\theta}:=(\theta_{i_{1}},\ldots,\theta_{i_{k}})$. Then $S$ is parametrized by the map
$V\ni\underline{\theta}\mapsto(\theta_{1}(\underline{\theta}),\ldots,\theta_{r}(\underline{\theta}))\in S\subset E.$
@@
$=\prod_{i=1}^{r}k_{i}(n_{i})\cdot\prod_{l=1}^{k}\binom{n_{i_{l}}+\alpha_{i_{l}}}{n_{i_{l}}}^{-1}P_{n_{i_{l}}}^{(\alpha_{i_{l}},\beta_{i_{l}})}(\cos(\theta_{i_{l}}-\theta_{i_{l}}^{\prime})),$
@@
$\mathcal{A}:g\mapsto\int_{V}g(\underline{\theta}^{\prime})\widetilde{\mathscr{K}}(\underline{\theta}^{\prime},\underline{\theta})\ d\underline{\theta}^{\prime}.$
Furthermore, extending $g$ by zero outside of $V$ in $\mathbb{T}^{k}:=\mathbb{T}_{i_{1}}\times\cdots\times\mathbb{T}_{i_{k}}$, it suffices to get the same bound for the $L^{p^{\prime}}(\mathbb{T}^{k},\ d\underline{\theta})\to L^{p}(\mathbb{T}^{k},\ d\underline{\theta})$ norm of the operator
$\widetilde{\mathcal{A}}:g\mapsto\int_{\mathbb{T}^{k}}g(\underline{\theta}^{\prime})\widetilde{\mathscr{K}}(\underline{\theta}^{\prime},\underline{\theta})\ d\underline{\theta}^{\prime},$
as $\|\mathcal{A}\|_{L^{p^{\prime}}(V,\ d\underline{\theta})\to L^{p}(V,\ d\underline{\theta})}\leq\|\widetilde{\mathcal{A}}\|_{L^{p^{\prime}}(\mathbb{T}^{k},\ d\underline{\theta})\to L^{p}(\mathbb{T}^{k},\ d\underline{\theta})}$. Then the above operator is convolution on $\mathbb{T}^{k}$ with a product of Jacobi polynomials. First observe that by the standard asymptotics
@@
and Lemma 2.2, we have
(2.11) $\prod_{i=1}^{r}k_{i}(n_{i})\cdot\prod_{l=1}^{k}\binom{n_{i_{l}}+\alpha_{i_{l}}}{n_{i_{l}}}^{-1}$ $\lesssim\prod_{i=1}^{r}(n_{i}+1)^{d_{i}-1}\prod_{l=1}^{k}(n_{i_{l}}+1)^{-\alpha_{i_{l}}}.$
Then noting that $p\geq p^{\prime}$ as $p\geq 2$, we apply the Minkowski inequality along with Theorem 1.8, to obtain
$\left\|f*\prod_{l=1}^{k}P_{n_{i_{l}}}^{(\alpha_{i_{l}},\beta_{i_{l}})}(\cos(\theta_{i_{l}}))\right\|_{L^{p}(\mathbb{T}^{k})}$ $\lesssim\prod_{l=1}^{k}\|T_{n_{i_{l}}}^{\alpha_{i_{l}},\beta_{i_{l}}}\|_{L^{p^{\prime}}(\mathbb{T})\to L^{p}(\mathbb{T})}\|f\|_{L^{p^{\prime}}(\mathbb{T}^{k})}$
(2.12) $\lesssim\prod_{l=1}^{k}(n_{i_{l}}+1)^{\alpha_{i_{l}}+2\tau(d_{i_{l}},p)}\|f\|_{L^{p^{\prime}}(\mathbb{T}^{k})}.$
Here in the last inequality we used $\alpha_{i_{l}}=(d_{i_{l}}-2)/2$. Combining (2.11) and (2.2) as well as the facts that $\tau(d,p)\geq-(d-1)/{4}$ and $n_{i}\lesssim N$, we conclude that
$\|\widetilde{\mathcal{A}}\|_{L^{p^{\prime}}(\mathbb{T}^{k})\to L^{p}(\mathbb{T}^{k})}\lesssim N^{{d-r}+2\sum_{l=1}^{k}\tau(d_{i_{l}},p)}\leq N^{{d-r}+2\sum_{i=1}^{k}\tau(d_{i},p)}.$
@@
Suppose $\Delta f=-N^{2}f$, $N>1$. By (2.4), and a standard counting estimate [3], we get
@@
We decompose the convolution kernel $P^{\alpha,\beta}_{n}(\cos\theta)$ according to the value of $\theta$, and it suffices to restrict attention to $\theta\in[0,\pi]$. Let $c$ be a fixed positive constant. Denote
$\widetilde{n}:=n+\frac{\alpha+\beta+1}{2}$
and
$\gamma:=-\frac{\pi}{2}\left(\alpha+\frac{1}{2}\right).$
Case 1. $c(n+1)^{-1}\leq\theta\leq\pi-c(n+1)^{-1}$. We use (8.21.18) of [11] as follows:
@@
Case 2. $0<\theta\leq c(n+1)^{-1}$ or $\pi-c(n+1)^{-1}\leq\theta<\pi$. For the former, we use (8.21.17) of [11] as follows:
@@
Here $J_{\alpha}$ stands for the Bessel function of the first kind of order $\alpha$, as defined in (1.71.1) of [11]. By (1.71.10) of [11], we have the asymptotics,
@@
A similar approach works for the other case $\pi-c(n+1)^{-1}\leq\theta<\pi$: combining (8.21.17) and (4.1.3) of [11], we get
@@
at the kink point $p={4}/{(2\alpha+1)}>2$. To accomplish this, we adjust the proof of Proposition 3.1: the only change needed is to apply the Hardy–Littlewood–Sobolev inequality in place of Young’s convolution inequality to (3.1) over the interval $c(n+1)^{-1},\pi/2]$, as well as over the interval $\pi/2,\pi-c(n+1)^{-1}]$ if $\beta=\alpha$; if $\beta<\alpha$, the original application of Young’s inequality over the interval $\pi/2,\pi-c(n+1)^{-1}]$ via (3.2) remains sufficient. For the remaining case $\alpha=\beta=1/2$, we use the explicit formula (4.1.7) of [11],
@@
As the spherical functions $\Phi_{i,n_{i}}$ are matrix entries of the associated irreducible spherical representations expressed via unit vectors (see Theorem 4.3 of Chapter V of [4]), the Schur orthogonality relations yield orthogonality between $\Phi_{i,n_{i}}$ for different $n_{i}$, and
@@
When $r\geq 5$, by the equidistribution of lattice points on $r$-dimensional spheres as established by Pommerenke [8], it holds $|\Lambda|\gtrsim N^{r-2}$ whenever $\Lambda\neq\varnothing$, which yields the sharpness of (1.4).